# Overcoming the Thermal-Noise Limit of Microwave Measurements by Pre-cooling with an Active Cold Load

### v2

@@ I Introduction
Though currently still a few factors of ten greater in temperature/photon number than what can be achieved through standard mK refrigeration (as involves the removal of thermal phonons from the materials that surround or fill an electromagnetic cavity), recent studiesWu et al. (2021); Ng et al. (2021); Fahey et al. (2023); Wang et al. (2024); Day et al. (2024) have demonstrated the direct removal of thermal photons from particular microwave modes inside bench-top cavities at room temperature through the stimulated absorption of these photons by an optically spin-polarised medium placed inside the cavity. While obviating the need for physical refrigeration, this hybrid (optical-microwave) approach still requires a sufficiently powerful optical pump source for spin-polarizing the absorbing medium. This results in still quite a bulky and power-hungry system, even when a miniaturized microwave cavity of high magnetic Purcell factorBreeze et al. (2015); Fleury et al. (2023) is used to reduce the optical pump power/energy needed.
@@
Within microwave instrumentation, electrically powered low-noise amplifiers (LNAs) are ubiquitously used to increase the amplitude (and thus power) of weak signals. Guided quantitatively by Friis’ formulaFriis (1944), the deployment of the quietest LNA available at the front end of an instrument’s amplification chain will minimize the amount of deleterious noise added to the measured signal, thereby maximize the overall signal-to-noise ratio (SNR) and measurement sensitivityŠimėnas et al. (2021). Compared to their manifest utility in radio-astronomyGiordmaine et al. (1959) and deep-space communicationClauss and Shell (2008), the gains in sensitivity from installing a lower-noise (cryogenic) preamplifier into an electron paramagnetic resonance (EPR) or nuclear magnetic resonance (NMR) spectrometer are found, empirically, to be both (i) modestCollier et al. (1968); Narkowicz et al. (2013); Kalendra et al. (2023), especially when the spectrometer’s microwave or radio-frequency (RF) cavity is not itself refrigerated, and (ii) dependent on the degree of cavity-amplifier couplingPfenninger et al. (1995).
@@
Various schemes such as Townes’ “Q-multiplier”Townes (1960), involving the incorporation of a maser gain medium inside the cavity itself, have been proposed over the years towards improving measurement sensitivity. It was Mollier et al.Mollier et al. (1973) who cut through the confusion to affirm that, no matter how low-noise the receiving LNA gain medium, the sensitivity of an EPR (or NMR) spectrometer is ultimately limited by thermal noise (i.e. incoherent microwave photons) generated from within the EPR cavity itself. An ultra-low-noise receiving amplifier attached directly to the cavity’s coupling port, or even an ultra-low-noise gain medium resident inside the cavity, will amplify both the signal and this noise equally; thus, the SNR won’t improve. As addressed quantitatively by PoolePoole and Farach (2019) and othersHill and Richards (1968); Rinard et al. (1999) since, the only ways to improve sensitivity (for a given sample, completely filling the cavity, at a given temperature) are to either increase the cavity’s quality factor $Q$ (albeit at the expense of measurement bandwidth), so increasing the photon cavity lifetime, or to increase the EPR transition frequency (by applying a strong d.c. magnetic field to the sample), so reducing the number of thermal photons in the cavity’s operational mode. Or else (back to where this argument began:) to physically cool the cavity or otherwise “state-prepare” its operational microwave mode(s).
@@
At least for frequencies in the low-GHz region, today’s solid-state LNAs, in the form of monolithic microwave integrated circuits (MMICs) or else as hybrids composed of interconnected discrete components, containing either InP- or GaAs-based high-electron-mobility transistors (HEMTs) or else SiGe-based heterojunction bipolar transistors (HBTs), offer impressively low noise temperatures (referred to input) of 50 K or less –even when operating at room temperature. As mature commercial-off-the-shelf (COTS) components, they are readily purchasable. If suitably designed/chosen, the noise temperature of such an LNA will decrease by a further order of magnitude or so upon cooling to liquid-helium temperatureWeinreb et al. (1982); Pospieszalski (2024). With such low amplifier noise temperatures available, the sensitivity with which the field-amplitude/energy ($=$ number of photons) inside a room-temperature microwave cavity can be measured is determined almost entirely by the cavity’s own thermal noise (as is associated with the rate at which electromagnetic energy is dissipated inside the cavity) –not the small amount of extra noise imparted by the receiving amplifier chain used to monitor the cavity’s field amplitude through a port attached to the cavity.
@@
As already mentioned above, Wu et al.Wu et al. (2021) demonstrated the cooling of an electromagnetic (EM) mode by coupling it to a “spin-cold” reservoir in the form of an optically excited crystal of pentacene-doped para-terphenyl (Pc:PTP). Ng et al.Ng et al. (2021), Blank et al.Blank et al. (2023), and Day et al.Day et al. (2024), have since demonstrated quasi-continuous mode cooling using photoexcited negatively charged nitrogen vacancies (NV-) in diamond as a spin-polarised absorber. In a similar vein, Englund, Trusheim and co-workersFahey et al. (2023); Wang et al. (2024) have exploited mode cooling with NV- diamond to enhance the sensitivity of a magnetometer. All of these works use a cm-sized “3-D” dielectrically-loaded cavity to support a TE01δ mode affording a high magnetic Purcell factorBreeze et al. (2015). The frequency $f_{0}$ of this mode is adjusted, or a d.c. magnetic field is applied, to align it with the absorptive paramagnetic transition being targeted. Since the linewidth of this transition is narrow (a few MHz for pentacene, a few hundreds of kHz for NV- diamond), these approaches to mode cooling are intrinsically narrow-band.
@@ II Active Cold Loads
In microwave radiometry, cooled coaxial or waveguide terminations, also known as “cold loads”, are used both as sources of noise and as absorbersGervasi et al. (1995) of unwanted signals. Such a cryogenic load is a passive component requiring no source of power. To be useful as an accurately quantified noise source or “reference”, it must be anchored thermally to a heat bath (such as a “cold plate” within a cryostat) whose temperature is accurately known. A cryogenic load can be connected to one of the ports of a (cryogenic) circulator to form an isolator. At least within the circulator’s operational bandwidth, this isolator can be used to protect a sensitive quantum circuit from backwardly propagating thermal noise and/or coherent signals unintentionally generated by reflections further up/along the signal path. In this context, we point out that the pre-cooling method introduced here draws parallels with the radiative cooling of a spin ensemble or resonator, held physically at liquid-helium temperature, by coupling it electromagnetically to an absorber (in the form of either a passive load or active device) held at mK temperatureAlbanese et al. (2020); Wang et al. (2021).
@@
A convenient alternative to a physically cryogenic termination is an active cryogenic load (ACL)Frater and Williams (1981); Forward and Terry (2010), which, though a powered device, mimics the former’s behavior. ACLs can be made from the inputs (gates/bases) of suitably biased low-noise transistors configured as single-port terminated amplifiersDunleavy et al. (1997); Leynia de la Jarrige et al. (2010); Weissbrodt (2017). Though not always cold enough (nor broad-band or stable enough) to replace truly cryogenic terminations in all applications, they can be used advantageously in airborne and space applications to calibrate (or at least drift-correct) earth-monitoring radiometersShou et al. (2015); Weissbrodt (2017), dispensing with the encumbrance of mobile refrigeration and/or the operational need to “roll” the instrument’s platform to point the radiometer at dark sky (i.e., at 2.7 K cosmic background radiation –serving as an absolute temperature reference).
Acknowledging that other approaches existClerk et al. (2010), we here apply the so-called “wave approach” to analyse the generation and flow of noise in realistic microwave circuits (see chapter 8 of Ref. Siegman (1964), or Refs. Meys (1978); Dorđević et al. (2015)). Through the construction of flow graphs combined with the application of Mason’s “non-touching loop” ruleHewlett Packard / Agilent (1990); Hecken (1981), this diagrammatic approach readily encompasses the frequently encountered complexity of what a laser physicist would describe as “etalon effects”, where simultaneously forward- and back-propagating noise waves form (partial) standing waves along sections of transmission line between two or more impedance mismatches (the equivalent of partially reflecting mirrors functioning as beamsplitters/combiners). The available noise temperature from the input of a low-noise amplifier (of sufficiently high gain) that is terminated by a matched and sufficiently cold load can thereby be quantified asMeys (1978); Escotte (2015):
@@
Note that the minus sign in front of $T_{\text{min}}$ in Eq. 1 can lead to (partial) cancellation between the two terms on the r.h.s. of the same, all depending on the exact magnitude of $R_{\text{n}}$ and the locations of $S_{11}$ and $\Gamma_{\text{opt}}$ in the complex plane. For the particular LNA used in our room-temperature experiments reported here, namely a Qorvo QPL9547, we calculate, from its specification sheetQorvo (2023), that $T_{\text{cold}}~{}=~{}29.1$ K at 1.45 GHz –around 10 times lower than the LNA’s operating temperature.
Alas, complete sets of scattering and noise parameters are rarely published for cryogenic amplifiers, preventing immediate off-paper predictions of their performance when repurposed as cryo-ACLs. One informative exception is the noise calibration, exploiting a cryogenic impedance tuner, of the type of LNA used for the “ALPACA” phased-array radio telescope, as reported in ref. Sheldon et al. (2021). Though only the magnitude of $S_{11}$ (not its angle) is provided, upon reading off from Figs. 12 and 13, ibid., one can frame this LNA’s available input noise temperature to within: 6.2 K $<T_{\text{rev}}<$ 7.3 K at 1.5 GHz, when operating at 20 K. Equivalent characterizations (including complex scattering parameters) of transistor-based, maser, and parametric amplifiers operating at even lower temperatures would enable the rational design of improved cryo-ACLs.
@@
Finally, it is instructive to compare Eq. 1 against the standard expression for the effective noise temperature of an LNAWait and Engen (1991); Dunleavy (2009); Modelithics + Mini-Circuits (2015) when used, as intended, as an amplifier:
@@
where $\Gamma_{\text{s}}$ is the coefficient of reflection off the source connected to the input of the LNA –in our case the port of a microwave cavity (see Fig. 1), possibly with a matching circuit/stub interposed. This expression will be needed later on for determining the temperature of a microwave mode –see Supplemental Materialsup (2024). $T_{\text{eff.}}$ can be minimized by “noise matching”, i.e. adjusting $\Gamma_{\text{s}}$ to equal $\Gamma_{\text{opt}}$ in the complex plane (viz. on the Smith chart). Note that minimizing $T_{\text{cold}}$ for an active cold load constructed from an LNA is a different task (their respective formulae 1 and 2 above are different).
@@ III Active Pre-Cooling
The first, cooling step is depicted in Figs. 1 and 2(a). In the bench-top experiments we report in this article, the physical walls and contents of the cavity, and the ACL attached to it, all lie at room temperature (nominally, 290 K). Note that the ACL is an electrically powered, single-port device. The output from the LNA that realizes the ACL does not (need to) go anywhere; it can be sunk into a matched passive load at ambient temperature or, if the attenuation backwards through the cooling LNA (essentially $|S12|$) is insufficient, into a second ACL –provided by the input of an additional LNA whose output is terminated by an ambient matched load. [We point out parenthetically that the idea of advantageously terminating the output of the “primary” cooling LNA with a “secondary” cooling LNA (and so on in cascade, as far as needs be) is not newLardizabal (1997).] In our experiments, the input of the cooling LNA is connected via a low-insertion-loss switch to a large loop of wire located inside the cavity, which serves as an over-coupled port.
@@
For understanding and quantifying the effects of coupling an external ACL to a microwave cavity, we now appeal to what Siegman describes simply as a “useful noise theorem” in section 8.3 of his bookSiegman (1964). Consider a cavity whose internal electromagnetic field is monitored by a small (weakly coupled) probe attached to an external port, so providing an output (“O/P”; see Fig. 3 –the equivalent of Fig. 8-10, ibid.). Here, one should regard the cavity’s field as the common “bus” (or “loss-less network” –in Siegman’s more circuit-oriented parlance) through which all of it components, including external loads, interact. Applying the theorem: The available noise power per unit bandwidth $B$ of an electromagnetic mode inside the cavity expressed in units of temperature, namely $T_{\text{mode}}=P_{\text{O/P}}/[k_{\text{B}}B(1-|\Gamma_{\text{O/P}}|^{2})]$, will equal the weighted sum of the temperatures $T_{i}$ of the $N$ bodies that dissipate the mode’s energy, where the weighting $p_{i}$ given to each body equals the fraction of the mode’s energy that it dissipates, and where the sum of these fractions adds up to unity: $\sum_{i}^{N}p_{i}=1$. Mathematically (cf. equation 8-3-3 in Ref. Siegman (1964)): $T_{\text{mode}}=\sum_{i}^{N}p_{i}T_{i}$. We stress that $T_{\text{mode}}$ quantifies the available noise power (per unit bandwidth), taking into account, through the above factor of $(1-|\Gamma_{\text{O/P}}|^{2})]$, the rate at which noise energy is extracted by the probe, where $\Gamma_{\text{O/P}}$ is the reflection coefficient off the probe’s port (again, see Fig. 3). For the sake of the argument here, we assume that the probe’s coupling and thus the noise power the probe itself extracts, as will get sunk into a matched load (at the input of a monitoring instrument), is infinitessimally small, such that this load’s dissipation does not contribute to the $\sum_{i}^{N}$ sum. One must regard the cavity’s enclosing metal walls (of finite conductivity), any dielectric or magnetic material (of finite loss tangent) inside the cavity, plus any other components that contribute to the mode’s intrinsic loss, as such dissipating bodies. External loads coupled to the cavity, such as an active cryogenic load, also count. The thermodynamic temperature that the electromagnetic mode itself adopts is a dissipation-weighted balance between them all.
@@
One immediately sees from this equation that, for a high coupling factor $\kappa=\kappa_{\text{over}}\gg 1$, corresponding to a substantially over-coupled port, the microwave cavity’s noise temperature will approach that of the cold load, $T_{\text{cold}}$. The dynamics of the cooling process is treated around Eqs. S1 and S2 in the Supplemental Materialsup (2024), whose steady-state solution (S3) agrees with the thermodynamic treatment above. The time scale on which the mode is cooled is set by the mode’s as-loaded photon lifetime, namely $Q_{0}/[2\pi f_{\text{mode}}(1+\kappa_{\text{over}})]$.
@@
In the case of an active cold load formed by the (gated) channel of the LNA’s input HEMT, the height of this channel is just a few nm (defining a layer of two-dimensional electron gas: a “2-DEG”) and its lateral dimensions (length and width) are sub-micronsBaustista (2008). The cooling capacity of this atto-liter volume of semiconductor is limited. Operationally, one requires that the density, in frequency space, of the electromagnetic modes to be cooled is sufficiently low such that the noise power provided by the total number of modes within the input HEMT’s operating bandwidth is insufficient to saturate/overload its channel. This means that the cavity enclosing and defining the modes must not be too large (to prevent the spectral density of modes becoming too high) nor too hot, and the coupling factors of its modes to the ACL must not be too high. As a last resort, saturation could be thwarted by inserting a low-loss band-pass filter between the cooling LNA and the cavity to act as a pre-selector, so coupling the ACL to fewer modes. A sense of scale is useful with regards to these concerns, however: The noise power impinging on the LNA’s input, at a coupling factor of $\kappa$, from a cavity mode at absolute temperature $T$, equals $2\pi f_{\text{mode}}\kappa k_{\text{B}}T/Q_{0},$ where $k_{\text{B}}$ is Boltzmann’s constant. For the experimental parameters of the demonstrations reported in this article, this amounts to -121 dBm per mode. The combined power from the several dozen modes lying within our LNA’s frequency range (for the dimensions and dielectric loading of the cavity that we used) is thus still around ten orders of magnitude weaker than our LNA’s specified saturation input power (at -1 dB compression). How the LNA’s noise parameters themselves depend on the input (noise) power at elevated power levels is seldom reported, though this huge ratio suggests that the cooling capacity of an ACL constructed from a solid-state LNA enjoys substantial headroom.
@@
Figure 4: (a) Photograph of the experimental resonator’s interior: a sapphire ring holding a 0.1% Pc:PTP crystal, surrounded by the silver-plated walls of a cylindrical brass enclosure. (b) CAD diagram of the same resonator with a two-dimensional axisymmetric simulation of the magnetic flux (white arrows) and magnetic energy density of the TE01δ microwave modeWu et al. (2020a) that this resonator supports. (c) Signal-path anatomy of the experimental time-resolved EPR spectrometer used to demonstrate active pre-cooling. This circuit incorporates homodyne detection of the transient EPR signal. The firing of the optical-parametric oscillator (OPO) laser system and the closing/opening of the two microwave switches are scheduled via transistor–transistor logic (TTL) control pulses (generated by two cascaded pulse generators –both a Thandar TG105) as explained in Fig. 7.
@@
[2] The noise temperature on the same port as seen/felt by the cavity will also be slightly higher than the cooling LNA’s (available) input noise temperature due to insertion losses (I.L.) along this same connection. These losses comprise [see Fig. 4(c)] the I.L. of the single-pole double-throw (SPDT) switch itself plus those of the two short connecting coaxial cables on either side of it. The increased noise temperature generated by them equals $\lambda T_{\text{0}}$, with $\lambda\equiv L_{\text{dB}}/4.34$, $L_{\text{dB}}$ being the overall I.L. in units of dBSiegman (1964); we measured this loss to be 0.19 dB.
@@
Including all of the above additional contributions (as per Eq. S13 of the Supplemental Materialsup (2024)), the lowest obtainable mode temperature, $T_{\text{mode}}^{\text{cooled}}$, is predicted to be
@@
where $T_{\text{mon.}}$ is the effective cooling temperature of the monitoring LNA taking into account the unavoidable I.L. of the triple-stub tuner connecting it to the cavity’s monitoring port; this was measured to be 6.05 dB and takes $T_{\text{mon.}}$ up to 225.2 K. The expected noise temperature inside the cavity after pre-cooling, as calculated from Eq. 4 above (see Eq. S16 in the Supplemental Materialsup (2024)), is $T_{\text{mode}}^{\text{cooled}}$ = 131.5 K.

### v3

@@ I Introduction
Though currently still a few factors of ten greater in temperature/photon number than what can be achieved through standard mK refrigeration (as involves the removal of thermal phonons from the materials that surround or fill an electromagnetic cavity), recent studies[1, 2, 3, 4, 5] have demonstrated the direct removal of thermal photons from particular microwave modes inside bench-top cavities at room temperature through the stimulated absorption of these photons by an optically spin-polarised medium placed inside the cavity. While obviating the need for physical refrigeration, this hybrid (optical-microwave) approach still requires a sufficiently powerful optical pump source for spin-polarizing the absorbing medium. This results in still quite a bulky and power-hungry system, even when a miniaturized microwave cavity of high magnetic Purcell factor[6, 7] is used to reduce the optical pump power/energy needed.
@@
Within microwave instrumentation, electrically powered low-noise amplifiers (LNAs) are ubiquitously used to increase the amplitude (and thus power) of weak signals. Guided quantitatively by Friis’ formula[8], the deployment of the quietest LNA available at the front end of an instrument’s amplification chain will minimize the amount of deleterious noise added to the measured signal, thereby maximize the overall signal-to-noise ratio (SNR) and measurement sensitivity[9]. Compared to their manifest utility in radio-astronomy[10] and deep-space communication[11], the gains in sensitivity from installing a lower-noise (cryogenic) preamplifier into an electron paramagnetic resonance (EPR) or nuclear magnetic resonance (NMR) spectrometer are found, empirically, to be both (i) modest[12, 13, 14], especially when the spectrometer’s microwave or radio-frequency (RF) cavity is not itself refrigerated, and (ii) dependent on the degree of cavity-amplifier coupling[15].
@@
Various schemes such as Townes’ “Q-multiplier”[16], involving the incorporation of a maser gain medium inside the cavity itself, have been proposed over the years towards improving measurement sensitivity. It was Mollier et al.[17] who cut through the confusion to affirm that, no matter how low-noise the receiving LNA gain medium, the sensitivity of an EPR (or NMR) spectrometer is ultimately limited by thermal noise (i.e. incoherent microwave photons) generated from within the EPR cavity itself. An ultra-low-noise receiving amplifier attached directly to the cavity’s coupling port, or even an ultra-low-noise gain medium resident inside the cavity, will amplify both the signal and this noise equally; thus, the SNR won’t improve. As addressed quantitatively by Poole[18] and others[19, 20] since, the only ways to improve sensitivity (for a given sample, completely filling the cavity, at a given temperature) are to either increase the cavity’s quality factor $Q$ (albeit at the expense of measurement bandwidth), so increasing the photon cavity lifetime, or to increase the EPR transition frequency (by applying a strong d.c. magnetic field to the sample), so reducing the number of thermal photons in the cavity’s operational mode. Or else (back to where this argument began:) to physically cool the cavity or otherwise “state-prepare” its operational microwave mode(s).
@@
At least for frequencies in the low-GHz region, today’s solid-state LNAs, in the form of monolithic microwave integrated circuits (MMICs) or else as hybrids composed of interconnected discrete components, containing either InP- or GaAs-based high-electron-mobility transistors (HEMTs) or else SiGe-based heterojunction bipolar transistors (HBTs), offer impressively low noise temperatures (referred to input) of 50 K or less –even when operating at room temperature. As mature commercial-off-the-shelf (COTS) components, they are readily purchasable. If suitably designed/chosen, the noise temperature of such an LNA will decrease by a further order of magnitude or so upon cooling to liquid-helium temperature[21, 22]. With such low amplifier noise temperatures available, the sensitivity with which the field-amplitude/energy ($=$ number of photons) inside a room-temperature microwave cavity can be measured is determined almost entirely by the cavity’s own thermal noise (as is associated with the rate at which electromagnetic energy is dissipated inside the cavity) –not the small amount of extra noise imparted by the receiving amplifier chain used to monitor the cavity’s field amplitude through a port attached to the cavity.
@@
As already mentioned above, Wu et al.[1] demonstrated the cooling of an electromagnetic (EM) mode by coupling it to a “spin-cold” reservoir in the form of an optically excited crystal of pentacene-doped para-terphenyl (Pc:PTP). Ng et al.[2], Blank et al.[23], and Day et al.[5], have since demonstrated quasi-continuous mode cooling using photoexcited negatively charged nitrogen vacancies (NV-) in diamond as a spin-polarised absorber. In a similar vein, Englund, Trusheim and co-workers[3, 4] have exploited mode cooling with NV- diamond to enhance the sensitivity of a magnetometer. All of these works use a cm-sized “3-D” dielectrically-loaded cavity to support a TE01δ mode affording a high magnetic Purcell factor[6]. The frequency $f_{0}$ of this mode is adjusted, or a d.c. magnetic field is applied, to align it with the absorptive paramagnetic transition being targeted. Since the linewidth of this transition is narrow (a few MHz for pentacene, a few hundreds of kHz for NV- diamond), these approaches to mode cooling are intrinsically narrow-band.
@@ II Active Cold Loads
In microwave radiometry, cooled coaxial or waveguide terminations, also known as “cold loads”, are used both as sources of noise and as absorbers[24] of unwanted signals. Such a cryogenic load is a passive component requiring no source of power. To be useful as an accurately quantified noise source or “reference”, it must be anchored thermally to a heat bath (such as a “cold plate” within a cryostat) whose temperature is accurately known. A cryogenic load can be connected to one of the ports of a (cryogenic) circulator to form an isolator. At least within the circulator’s operational bandwidth, this isolator can be used to protect a sensitive quantum circuit from backwardly propagating thermal noise and/or coherent signals unintentionally generated by reflections further up/along the signal path. In this context, we point out that the pre-cooling method introduced here draws parallels with the radiative cooling of a spin ensemble or resonator, held physically at liquid-helium temperature, by coupling it electromagnetically to an absorber (in the form of either a passive load or active device) held at mK temperature[25, 26].
@@
A convenient alternative to a physically cryogenic termination is an active cryogenic load (ACL)[27, 28], which, though a powered device, mimics the former’s behavior. ACLs can be made from the inputs (gates/bases) of suitably biased low-noise transistors configured as single-port terminated amplifiers[29, 30, 31]. Though not always cold enough (nor broad-band or stable enough) to replace truly cryogenic terminations in all applications, they can be used advantageously in airborne and space applications to calibrate (or at least drift-correct) earth-monitoring radiometers[32, 31], dispensing with the encumbrance of mobile refrigeration and/or the operational need to “roll” the instrument’s platform to point the radiometer at dark sky (i.e., at 2.7 K cosmic background radiation –serving as an absolute temperature reference).
Acknowledging that other approaches exist[33], we here apply the so-called “wave approach” to analyse the generation and flow of noise in realistic microwave circuits (see chapter 8 of Ref. [34], or Refs. [35, 36]). Through the construction of flow graphs combined with the application of Mason’s “non-touching loop” rule[37, 38], this diagrammatic approach readily encompasses the frequently encountered complexity of what a laser physicist would describe as “etalon effects”, where simultaneously forward- and back-propagating noise waves form (partial) standing waves along sections of transmission line between two or more impedance mismatches (the equivalent of partially reflecting mirrors functioning as beamsplitters/combiners). The available noise temperature from the input of a low-noise amplifier (of sufficiently high gain) that is terminated by a matched and sufficiently cold load can thereby be quantified as[35, 39]:
@@
Note that the minus sign in front of $T_{\text{min}}$ in Eq. 1 can lead to (partial) cancellation between the two terms on the r.h.s. of the same, all depending on the exact magnitude of $R_{\text{n}}$ and the locations of $S_{11}$ and $\Gamma_{\text{opt}}$ in the complex plane. For the particular LNA used in our room-temperature experiments reported here, namely a Qorvo QPL9547, we calculate, from its specification sheet[44], that $T_{\text{cold}}~{}=~{}29.1$ K at 1.45 GHz –around 10 times lower than the LNA’s operating temperature.
Alas, complete sets of scattering and noise parameters are rarely published for cryogenic amplifiers, preventing immediate off-paper predictions of their performance when repurposed as cryo-ACLs. One informative exception is the noise calibration, exploiting a cryogenic impedance tuner, of the type of LNA used for the “ALPACA” phased-array radio telescope, as reported in ref. [45]. Though only the magnitude of $S_{11}$ (not its angle) is provided, upon reading off from Figs. 12 and 13, ibid., one can frame this LNA’s available input noise temperature to within: 6.2 K $<T_{\text{rev}}<$ 7.3 K at 1.5 GHz, when operating at 20 K. Equivalent characterizations (including complex scattering parameters) of transistor-based, maser, and parametric amplifiers operating at even lower temperatures would enable the rational design of improved cryo-ACLs.
@@
Finally, it is instructive to compare Eq. 1 against the standard expression for the effective noise temperature of an LNA[42, 40, 41] when used, as intended, as an amplifier:
@@
where $\Gamma_{\text{s}}$ is the coefficient of reflection off the source connected to the input of the LNA –in our case the port of a microwave cavity (see Fig. 1), possibly with a matching circuit/stub interposed. This expression will be needed later on for determining the temperature of a microwave mode –see Supplemental Material[46] (which includes references [47, 48, 49]). $T_{\text{eff.}}$ can be minimized by “noise matching”, i.e. adjusting $\Gamma_{\text{s}}$ to equal $\Gamma_{\text{opt}}$ in the complex plane (viz. on the Smith chart). Note that minimizing $T_{\text{cold}}$ for an active cold load constructed from an LNA is a different task (their respective formulae 1 and 2 above are different).
@@ III Active Pre-Cooling
The first, cooling step is depicted in Figs. 1 and 2(a). In the bench-top experiments we report in this article, the physical walls and contents of the cavity, and the ACL attached to it, all lie at room temperature (nominally, 290 K). Note that the ACL is an electrically powered, single-port device. The output from the LNA that realizes the ACL does not (need to) go anywhere; it can be sunk into a matched passive load at ambient temperature or, if the attenuation backwards through the cooling LNA (essentially $|S12|$) is insufficient, into a second ACL –provided by the input of an additional LNA whose output is terminated by an ambient matched load. [We point out parenthetically that the idea of advantageously terminating the output of the “primary” cooling LNA with a “secondary” cooling LNA (and so on in cascade, as far as needs be) is not new[50].] In our experiments, the input of the cooling LNA is connected via a low-insertion-loss switch to a large loop of wire located inside the cavity, which serves as an over-coupled port.
@@
For understanding and quantifying the effects of coupling an external ACL to a microwave cavity, we now appeal to what Siegman describes simply as a “useful noise theorem” in section 8.3 of his book[34]. Consider a cavity whose internal electromagnetic field is monitored by a small (weakly coupled) probe attached to an external port, so providing an output (“O/P”; see Fig. 3 –the equivalent of Fig. 8-10, ibid.). Here, one should regard the cavity’s field as the common “bus” (or “loss-less network” –in Siegman’s more circuit-oriented parlance) through which all of it components, including external loads, interact. Applying the theorem: The available noise power per unit bandwidth $B$ of an electromagnetic mode inside the cavity expressed in units of temperature, namely $T_{\text{mode}}=P_{\text{O/P}}/[k_{\text{B}}B(1-|\Gamma_{\text{O/P}}|^{2})]$, will equal the weighted sum of the temperatures $T_{i}$ of the $N$ bodies that dissipate the mode’s energy, where the weighting $p_{i}$ given to each body equals the fraction of the mode’s energy that it dissipates, and where the sum of these fractions adds up to unity: $\sum_{i}^{N}p_{i}=1$. Mathematically (cf. equation 8-3-3 in Ref. [34]): $T_{\text{mode}}=\sum_{i}^{N}p_{i}T_{i}$. We stress that $T_{\text{mode}}$ quantifies the available noise power (per unit bandwidth), taking into account, through the above factor of $(1-|\Gamma_{\text{O/P}}|^{2})]$, the rate at which noise energy is extracted by the probe, where $\Gamma_{\text{O/P}}$ is the reflection coefficient off the probe’s port (again, see Fig. 3). For the sake of the argument here, we assume that the probe’s coupling and thus the noise power the probe itself extracts, as will get sunk into a matched load (at the input of a monitoring instrument), is infinitessimally small, such that this load’s dissipation does not contribute to the $\sum_{i}^{N}$ sum. One must regard the cavity’s enclosing metal walls (of finite conductivity), any dielectric or magnetic material (of finite loss tangent) inside the cavity, plus any other components that contribute to the mode’s intrinsic loss, as such dissipating bodies. External loads coupled to the cavity, such as an active cryogenic load, also count. The thermodynamic temperature that the electromagnetic mode itself adopts is a dissipation-weighted balance between them all.
@@
One immediately sees from this equation that, for a high coupling factor $\kappa=\kappa_{\text{over}}\gg 1$, corresponding to a substantially over-coupled port, the microwave cavity’s noise temperature will approach that of the cold load, $T_{\text{cold}}$. The dynamics of the cooling process is treated around Eqs. S1 and S2 in the Supplemental Material[46], whose steady-state solution (S3) agrees with the thermodynamic treatment above. The time scale on which the mode is cooled is set by the mode’s as-loaded photon lifetime, namely $Q_{0}/[2\pi f_{\text{mode}}(1+\kappa_{\text{over}})]$.
@@
In the case of an active cold load formed by the (gated) channel of the LNA’s input HEMT, the height of this channel is just a few nm (defining a layer of two-dimensional electron gas: a “2-DEG”) and its lateral dimensions (length and width) are sub-microns[51]. The cooling capacity of this atto-liter volume of semiconductor is limited. Operationally, one requires that the density, in frequency space, of the electromagnetic modes to be cooled is sufficiently low such that the noise power provided by the total number of modes within the input HEMT’s operating bandwidth is insufficient to saturate/overload its channel. This means that the cavity enclosing and defining the modes must not be too large (to prevent the spectral density of modes becoming too high) nor too hot, and the coupling factors of its modes to the ACL must not be too high. As a last resort, saturation could be thwarted by inserting a low-loss band-pass filter between the cooling LNA and the cavity to act as a pre-selector, so coupling the ACL to fewer modes. A sense of scale is useful with regards to these concerns, however: The noise power impinging on the LNA’s input, at a coupling factor of $\kappa$, from a cavity mode at absolute temperature $T$, equals $2\pi f_{\text{mode}}\kappa k_{\text{B}}T/Q_{0},$ where $k_{\text{B}}$ is Boltzmann’s constant. For the experimental parameters of the demonstrations reported in this article, this amounts to -121 dBm per mode. The combined power from the several dozen modes lying within our LNA’s frequency range (for the dimensions and dielectric loading of the cavity that we used) is thus still around ten orders of magnitude weaker than our LNA’s specified saturation input power (at -1 dB compression). How the LNA’s noise parameters themselves depend on the input (noise) power at elevated power levels is seldom reported, though this huge ratio suggests that the cooling capacity of an ACL constructed from a solid-state LNA enjoys substantial headroom.
@@
Figure 4: (a) Photograph of the experimental resonator’s interior: a sapphire ring holding a 0.1% Pc:PTP crystal, surrounded by the silver-plated walls of a cylindrical brass enclosure. (b) CAD diagram of the same resonator with a two-dimensional axisymmetric simulation of the magnetic flux (white arrows) and magnetic energy density of the TE01δ microwave mode[52] that this resonator supports. (c) Signal-path anatomy of the experimental time-resolved EPR spectrometer used to demonstrate active pre-cooling. This circuit incorporates homodyne detection of the transient EPR signal. The firing of the optical-parametric oscillator (OPO) laser system and the closing/opening of the two microwave switches are scheduled via transistor–transistor logic (TTL) control pulses (generated by two cascaded pulse generators –both a Thandar TG105) as explained in Fig. 7.
@@
[2] The noise temperature on the same port as seen/felt by the cavity will also be slightly higher than the cooling LNA’s (available) input noise temperature due to insertion losses (I.L.) along this same connection. These losses comprise [see Fig. 4(c)] the I.L. of the single-pole double-throw (SPDT) switch itself plus those of the two short connecting coaxial cables on either side of it. The increased noise temperature generated by them equals $\lambda T_{\text{0}}$, with $\lambda\equiv L_{\text{dB}}/4.34$, $L_{\text{dB}}$ being the overall I.L. in units of dB[34]; we measured this loss to be 0.19 dB.
@@
Including all of the above additional contributions (as per Eq. S13 of the Supplemental Material[46]), the lowest obtainable mode temperature, $T_{\text{mode}}^{\text{cooled}}$, is predicted to be
@@
where $T_{\text{mon.}}$ is the effective cooling temperature of the monitoring LNA taking into account the unavoidable I.L. of the triple-stub tuner connecting it to the cavity’s monitoring port; this was measured to be 6.05 dB and takes $T_{\text{mon.}}$ up to 225.2 K. The expected noise temperature inside the cavity after pre-cooling, as calculated from Eq. 4 above (see Eq. S16 in the Supplemental Material[46]), is $T_{\text{mode}}^{\text{cooled}}$ = 131.5 K.